OpenStudy (anonymous):

Simplify. 16÷(5+3)·2 A.) 4 B.) 1 C.) 0 D.) ⁻1

OpenStudy (anonymous):

first add 5+3 then take that answer and multiply it by 2 then divide that answer by 16

OpenStudy (anonymous):

5+3=8 8*2=16 16/16=1

OpenStudy (anonymous):

Think its 4 coz after u used the Rule PEMDAS there is no separated so you will start again left to right .

OpenStudy (anonymous):

So the Answer is 4

OpenStudy (anonymous):

I forgot I was on here.. Lol thanks !

OpenStudy (anonymous):

Am I correct ?